OpenStudy (anonymous):

Find the slope of the line that passes through the points (2, 1) and (-1, -1).

OpenStudy (angela210793):

Slope=(y2-y1)/(x2-1)=-2/-3=2/3

OpenStudy (deez):

2/3

OpenStudy (anonymous):

\[ \Delta y \over \Delta x\]\[-1-1 \over -1 -2\]\[-2 \over -3\]\[2 \over 3\]